(b) If 3.0% of the heat being measured was being lost to the universe, determine whether they could they distinguish between iron (s = 0.444 J/g · °C) and Ni (s = 0.449 J/g · °C).

What is the percent difference between the specific heats of the two metals? Use the specific heat of iron as the theoretical value.
%

(b) Is this difference large enough to differentiate between the metals?

Yes

No   

This cannot be determined with the information given.

It depends on how hot the metal is before it is added it to the calorimeter.

It depends on the amount of water used in the calorimeter.
Answer #1

The percentage difference b/w the specific heats of the two metals are

= 0.449-0.444 / 0.444

= 0.01126 * 100 = 1.126 %

b)It depends on the amount of water used in the calorimeter.